# A Quarter of a Century of Neuromorphic Architectures on FPGAs - an Overview

### v2

@@ 2.2.2. Synapse Models
Neurons connect through synapses, which accept incoming spikes from other neurons. Synapses convert those spikes into a current or a change in conductance of the targeting neuron’s extensions (called dendrites). The current or change in conductance modifies the neuron’s membrane potential and can cause it to spike. The choice and functionality of a synapse model has a critical impact on the complexity of the hardware to a much greater extent than the choice of neuron model, mainly because there are many orders of magnitude more synapses than there are neurons(Azevedo et al., 2009). Synapses, which from the biological point of view include GABA, NMDA and AMPA ion channels(Ben-Ari et al., 1997), can be modeled as current-based, where they modify the injected current into the target membrane or conductance-based (CUBA), where they modify the conductance (COBA). As with neurons, synaptic dynamics are typically described using ODEs. Synapses are extremely important and are thought to play an even larger role than neurons in an SNN. This is because synapses facilitate learning in the change of their strength (conductance). This process is called synaptic plasticity(Citri and Malenka, 2008), and there are many different models available today that can be used to facilitate learning in an SNN. Most models are Hebbian(Gerstner, 2011) in nature, meaning that they relate the increase or decrease in strength to the simultaneous activity of pre- and post-synaptic neurons. One of the more popular examples is STDP(Song et al., 2000), which modifies the strength of synapses according to how close in time a synapse induces a spike in the target neuron. Supporting plasticity in the synapse model significantly increases the computational cost of a synapse. Apart from STDP, there are other important plasticity models, e.g., Bienenstock–Cooper–Munro (BCM) rule(Bienenstock et al., 1982) based on observations in homeostatic plasticity. There are also other methods, which can be classified as: (i) backpropagation-related (e.g., SpikeProp(Bohte et al., 2000), Backpropagation Through Time (BPTT)(Werbos, 1988)), (ii) direct-feedback-related (e-prop(Bellec et al., 2020), DECOLLE(Kaiser et al., 2020)) and (iii) creating an equivalent ANN, performing learning based on Stochastic Gradient Descent (SGD) or Backpropagation (BP) and transferring the resulting weights to an SNN.
2.3. Proposed taxonomy for Neuromorphic Architectures
@@
Schuman et al.(Schuman et al., 2022) suggest that for a system to be called neuromorphic, it must fulfill the following requirements: (i) it must compute using neurons and synapses (as opposed to instructions of a classical computer), (ii) it must communicate with spikes (as opposed to multi-bit values in a classical system), (iii) it is massively parallel (that is, it can compute all neurons/synapses fully in parallel), (iv) it collocates computation and memory (unlike von Neumann-based classical system), (v) it operates in an event-driven manner (the network state can be updated asynchronously, i.e., the neurons are not updated in sequence). Table 2 contrasts these properties with that of a classical (von Neurmann) computer system.
Figure 2. Proposed taxonomy - classes.
@@
We propose a taxonomy for hardware neuromorphic architectures (NMAs) that builds upon that of Schuman et al., but relaxes the constraints. We argue that while a neuromorphic system must be programmed using neurons and synapses and communicate with spikes (i–ii), the remaining requirements (iii-v) are, in fact, architectural design decisions that define traits and properties of the implementation of a hardware neuromorphic architecture (NMA). To summarize, we propose that a NMA must perform computations using neurons and synapses and communicate with spikes, but it may also be massively parallel, be asynchronous, and/or co-locate computations and memory. Ultimately, this results in a taxonomy consisting of eight classes, shown in Figure 2. The latter three optional characteristics will be referred to as Traits for the remainder of this paper. It is important to mention that the Classes are not synonymous to grades, i.e., there are no better or worse Classes. Instead, the classification is intended to compare NMAs and implementations similar in terms of hardware and relate them to each other, like Flynn’s taxonomy (Skillicorn, 1988). Every Class comes with its own advantages and disadvantages.
@@
Table 2. Comparison between von Neumann architectures and NMAs according to Schuman et al. Adapted from the Figure in (Schuman et al., 2022).
@@
3. Overview of the existing NMAs
@@
3.1. Class 0 (Traits: none)
Figure 3. Simplified diagram of a Class 0 NMA.
@@ 3.1. Class 0 (Traits: none)
Early implementations. Waldemark et al. (1998) (Waldemark et al., 1998) presented a small system consisting of one neuron and two synapses, a part of a bigger network realizing a classification task. The network utilized 19% of available logic resources on Altera Flex 10K. Jung et al. (2005) (Jung et al., 2005) presented an NMA, where a multi-master system depends on the efficient data switch to route the data between the specific modules. They built an architecture upon the SP2INN(Mehrtash et al., 2003) system concept and realizes a single neuron with eight fan-in connections with STDP functionality. Glackin et al. (2005) (Glackin et al., 2005b, a) presented an architecture for realizing large-scale networks on FPGAs. They chose a LIF neuron model with dynamics approximated with Euler integration with 0.125 ms timestep. They used 18b fixed-point number representation for weights and membranes. The use case was an SNN realizing angle translation. On the Xilinx Virtex-II Pro device, the authors implemented four PEs and 400 synaptic units running at 100 MHz, resulting in 4200 neurons and 1964200 synapses achieving timestep computation time 4327 times slower than real-time. Nuno-Maganda et al. (2009) (Nuno-Maganda et al., 2009) present an NMA for recall and learning with a multi-layer FF SNN. Architecture-wise, the system comprised several arrays of PEs and Learning Modules (LMs), connected to global data and control buses via two routers. The authors tested a system with two arrays of PEs and two LMs (16 neurons) on a Xilinx Virtex-II Pro device. They chose a 16b fixed-point number representation. Rice et al. (2009) (Rice et al., 2009) presented an implementation of an SNN comprised of 9000 IZH neurons for character recognition in images. Architecture-wise, the system comprised several parallel pipelined time-domain multiplexed (TDM) PEs. The representation was 16b fixed-point numbers for weights and membranes. The implementation was tested with an FF-FC SNN with two layers for recognizing characters in 96x96px bitmaps and 48 classes. The system was clocked at 199 MHz on Xilinx Virtex-4 and had 25 PEs.
@@
Early modern implementations. Podobas et al. (2017) (Podobas and Matsuoka, 2017) presented a scalable NMA relying on a single PE for neuron state update. The system supported IZH and HH dynamics, which were obtained through two approximation schemes: Euler integration with a timestep of 0.1 ms for IZH and Runge-Kutta (RK) with a timestep of 0.02 ms for HH. They used A2A networks of 14000 neurons to test the architecture on Altera Stratix-V. The synaptic connections were created from a common axon, which reduced the number of parallel computations related to spike event delivery, as well as alleviated the necessity to keep the information about the delayed spike for every connection separately. Mostafa et al. (2017) (Mostafa et al., 2017) presented an architecture for SNNs with sparse spiking activity for solving classification tasks. The architecture comprised eight PEs capable of processing 256 neurons. They chose an IF neuron model with membrane and current represented with signed 16b fixed-point number, and weights represented as signed 8b number. The implemented network had an FF-FC topology (610 neurons) and achieved an accuracy of 96.98% for the MNIST dataset. Galindo Sanchez et al. (2017) (Galindo Sanchez and Nunez-Yanez, 2017) presented a system for realizing Streaming SNN (S2NN) featuring switching off the parts of the implementation that are not needed. The design flow and the architecture assumed the implementation of FF-FC networks, layers of which could be ”switched off” when not in use. The design was implemented on a Xilinx Zynq-7000 device, where the integrated ARM A9 core controlled the power management. Kousanakis et al. (2017) (Kousanakis et al., 2017) presented a system with LIF neurons on a multi-FPGA system consisting of four Xilinx Virtex-6 devices. The architecture supported learning with a BCM learning rule, as well as random synaptic current generation. The system consisted of 40 PEs responsible for calculating six neurons each, with interconnectivity information and background synaptic noise stored in external memory. Humaidi et al. (2018) (Humaidi and Kadhim, 2018) compared an ANN and an SNN implementations for letter recognition on FPGA. The SNN implementation supported STDP for learning the patterns. The benchmark consisted of four different letters (5x3 pixels). The SNN was reported as more area-efficient and faster than ANN. The topology used was deduced to be a 48-neuron (IZH) WTA network.
@@
Modern implementations. Ahn et al. (2020) (Ahn, 2020) presented an architecture for realizing networks of HH neurons. The presented implementation simulated 12 million HH neurons with eight. The communication depended on the data being propagated along the axon bus from and to the PEs. The system stored duplicated information about the network state in every PE. The authors chose a neuron timestep of 0.04 ms, and the system achieved the frequency of 300 MHz on Xilinx Virtex UltraScale+. The number representation was a single-precision floating point. Ogaki et al. (2021) (Ogaki and Sato, 2021) presented an approach to realize HH-based SNNs with a fully dataflow-based approach [^2] Using the similar platform as in (Cheung et al., 2016) and (Cheung et al., 2012)).. The entire system was fully pipelined. The factor differentiating this architecture from the others was an explicit focus on representing the locality of biological neuronal connections - the neurons were connected only to their neighbors in the grid (2D array topology). They showed that their architecture running a 10000-neuron system (at 300 MHz) on an Alveo U200 FPGA could be up-to 7.5x faster than a state-of-the-art CPU. Zhou et al. (2021) (Zhou and Hu, 2021) presented an NMA, which relied on a set of TDM PEs capable of simulating 1024 neurons each. They utilized a weight-sharing concept, where 16 16b weights were shared throughout the synapses of the network. The architecture supported a LIF neuron model with a linear decay, instead of exponential. The system was implemented on a Xilinx Virtex-6 device and tested with the MNIST dataset with 64 PEs to realize an FF-FC network (2010 neurons), achieving an accuracy of 98.41% after 10 timesteps.
3.2. Class 1 (Traits: fully parallel)
Figure 4. Simplified diagram of a Class 1 NMA.
@@ 3.2. Class 1 (Traits: fully parallel)
Early implementations. Maya et al. (2000) (Maya et al., 2000) presented an architecture that comprised multiple PEs that computed a pulsed neuron - three in total - to solve the XOR problem. The weights were obtained through BP in MATLAB, and the structure was clocked at 90 MHz on Virtex VX50-6. Xicotencatl et al. (2003) (Xicotencatl and Arias-Estrada, 2003) presented a network of 1120 pulsed neurons with 1000 synapses. The architecture was implemented on Virtex-V2000 and was clocked at 35.6 MHz. The topology of the implemented network was a 2D array of neurons.
Early modern implementations. Moctezuma et al. (2013) (Moctezuma et al., 2013) presented an architecture for simulating ion-channel dynamics and membrane voltage changes of neurons defined by HH and Traub neuron models. The models utilized floating-point arithmetic, and exponential functions were implemented as LUTs. The system was using a MicroBlaze processor for configuration and supervision, as well as communication between subunits. The entire system operated at 100 MHz and allowed the simulation of six different configurations of soma, dendrite and synapse types from the aforementioned neuron models.
3.3. Class 2 (Traits: collocated computation and memory)
@@
Early implementations. Schrauwen et al. (2008) (Schrauwen et al., 2008) presented an LSM-based recognition system for spoken digits. The system relied on serialized synapse processing and serialized arithmetic. Architecture-wise, the system consisted of a set of PEs driven by a controller block. The neuronal dynamics of an SRM model were computed in a time-multiplexed manner, but we deduced that the PEs were not pipelined. The PEs operated at 100 MHz on Xilinx Virtex-4. The authors implemented a network of 1600 neurons as a use case, to process the incoming 16 kHz audio signal with five PEs.
Figure 5. Simplified diagram of a Class 2 NMA.
@@
Thomas et al. (2009) (Thomas and Luk, 2009) presented an NMA using the IZH neuron model. The system allowed for simulation of 1024 neurons and 1024000 synapses, running at 100x real-time speed on a Xilinx Virtex-5 device clocked at 133 MHz. The system achieved a throughput of 2.26 GFLOPs with double-precision floating-point number representation. Architecture-wise, the system consisted of a synaptic unit, implementing four synapse pipelines, an accumulator and synaptic weight RAMs and a neuron unit with a single neuron update pipeline. Wildie et al. (2009) (Wildie et al., 2009) presented an NMA that included a gap junction model on top of classic electrochemical synapses. They chose IZH as a neuron model for this system. Architecture-wise, the system consisted of eight PEs and could simulate a network of 1000 randomly-connected neurons. Every PE consisted of three pipelines - for neuron update, spike propagation and gap junctions dynamics. The system supported 24-bit fixed-point numerical representation for the parameters and conductance delay. It was clocked at 20 MHz and achieved 5.8x speed-up over real-time operation.

### v3

@@ 2.2.2 Synapse Models
Neurons connect through synapses, which accept incoming spikes from other neurons. Synapses convert those spikes into a current or a change in conductance of the targeting neuron’s extensions (called dendrites). The current or change in conductance modifies the neuron’s membrane potential and can cause it to spike. The choice and functionality of a synapse model has a critical impact on the complexity of the hardware to a much greater extent than the choice of neuron model, mainly because there are many orders of magnitude more synapses than there are neurons[58]. Synapses, which from the biological point of view include GABA, NMDA and AMPA ion channels[59], can be modeled as current-based, where they modify the injected current into the target membrane or conductance-based (CUBA), where they modify the conductance (COBA). As with neurons, synaptic dynamics are typically described using ODEs. Synapses are extremely important and are thought to play an even larger role than neurons in an SNN. This is because synapses facilitate learning in the change of their strength (conductance). This process is called synaptic plasticity[60], and there are many different models available today that can be used to facilitate learning in an SNN. Most models are Hebbian[61] in nature, meaning that they relate the increase or decrease in strength to the simultaneous activity of pre- and post-synaptic neurons. One of the more popular examples is STDP[23], which modifies the strength of synapses according to how close in time a synapse induces a spike in the target neuron. Supporting plasticity in the synapse model significantly increases the computational cost of a synapse. Apart from STDP, there are other important plasticity models, e.g., Bienenstock–Cooper–Munro (BCM) rule[62] based on observations in homeostatic plasticity. There are also other methods, which can be classified as: (i) backpropagation-related (e.g., SpikeProp[63], Backpropagation Through Time (BPTT)[64]), (ii) direct-feedback-related (e-prop[25], DECOLLE[65]) and (iii) creating an equivalent ANN, performing learning based on Stochastic Gradient Descent (SGD) or Backpropagation (BP) and transferring the resulting weights to an SNN.
2.3 Proposed taxonomy for Neuromorphic Architectures
@@
Schuman et al.[66] suggest that for a system to be called neuromorphic, it must fulfill the following requirements: (i) it must compute using neurons and synapses (as opposed to instructions of a classical computer), (ii) it must communicate with spikes (as opposed to multi-bit values in a classical system), (iii) it is massively parallel (that is, it can compute all neurons/synapses fully in parallel), (iv) it collocates computation and memory (unlike von Neumann-based classical system), (v) it operates in an event-driven manner (the network state can be updated asynchronously, i.e., the neurons are not updated in sequence). Table 2 contrasts these properties with that of a classical (von Neurmann) computer system.
Figure 2: Proposed taxonomy - classes.
@@
We propose a taxonomy for hardware neuromorphic architectures (NMAs) that builds upon that of Schuman et al., but relaxes the constraints. We argue that while a neuromorphic system must be programmed using neurons and synapses and communicate with spikes (i–ii), the remaining requirements (iii-v) are, in fact, architectural design decisions that define traits and properties of the implementation of a hardware neuromorphic architecture (NMA). To summarize, we propose that a NMA must perform computations using neurons and synapses and communicate with spikes, but it may also be massively parallel, be asynchronous, and/or co-locate computations and memory. Ultimately, this results in a taxonomy consisting of eight classes, shown in Figure 2. The latter three optional characteristics will be referred to as Traits for the remainder of this paper. It is important to mention that the Classes are not synonymous to grades, i.e., there are no better or worse Classes. Instead, the classification is intended to compare NMAs and implementations similar in terms of hardware and relate them to each other, like Flynn’s taxonomy [67]. Every Class comes with its own advantages and disadvantages.
@@
Table 2: Comparison between von Neumann architectures and NMAs according to Schuman et al. Adapted from the Figure in [66].
@@
3 Overview of the existing NMAs
@@
3.1 Class 0 (Traits: none)
Figure 3: Simplified diagram of a Class 0 NMA.
@@ 3.1 Class 0 (Traits: none)
Early implementations. Waldemark et al. (1998) [55] presented a small system consisting of one neuron and two synapses, a part of a bigger network realizing a classification task. The network utilized 19% of available logic resources on Altera Flex 10K. Jung et al. (2005) [68] presented an NMA, where a multi-master system depends on the efficient data switch to route the data between the specific modules. They built an architecture upon the SP2INN[69] system concept and realizes a single neuron with eight fan-in connections with STDP functionality. Glackin et al. (2005) [70, 71] presented an architecture for realizing large-scale networks on FPGAs. They chose a LIF neuron model with dynamics approximated with Euler integration with 0.125 ms timestep. They used 18b fixed-point number representation for weights and membranes. The use case was an SNN realizing angle translation. On the Xilinx Virtex-II Pro device, the authors implemented four PEs and 400 synaptic units running at 100 MHz, resulting in 4200 neurons and 1964200 synapses achieving timestep computation time 4327 times slower than real-time. Nuno-Maganda et al. (2009) [72] present an NMA for recall and learning with a multi-layer FF SNN. Architecture-wise, the system comprised several arrays of PEs and Learning Modules (LMs), connected to global data and control buses via two routers. The authors tested a system with two arrays of PEs and two LMs (16 neurons) on a Xilinx Virtex-II Pro device. They chose a 16b fixed-point number representation. Rice et al. (2009) [73] presented an implementation of an SNN comprised of 9000 IZH neurons for character recognition in images. Architecture-wise, the system comprised several parallel pipelined time-domain multiplexed (TDM) PEs. The representation was 16b fixed-point numbers for weights and membranes. The implementation was tested with an FF-FC SNN with two layers for recognizing characters in 96x96px bitmaps and 48 classes. The system was clocked at 199 MHz on Xilinx Virtex-4 and had 25 PEs.
@@
Early modern implementations. Podobas et al. (2017) [32] presented a scalable NMA relying on a single PE for neuron state update. The system supported IZH and HH dynamics, which were obtained through two approximation schemes: Euler integration with a timestep of 0.1 ms for IZH and Runge-Kutta (RK) with a timestep of 0.02 ms for HH. They used A2A networks of 14000 neurons to test the architecture on Altera Stratix-V. The synaptic connections were created from a common axon, which reduced the number of parallel computations related to spike event delivery, as well as alleviated the necessity to keep the information about the delayed spike for every connection separately. Mostafa et al. (2017) [74] presented an architecture for SNNs with sparse spiking activity for solving classification tasks. The architecture comprised eight PEs capable of processing 256 neurons. They chose an IF neuron model with membrane and current represented with signed 16b fixed-point number, and weights represented as signed 8b number. The implemented network had an FF-FC topology (610 neurons) and achieved an accuracy of 96.98% for the MNIST dataset. Galindo Sanchez et al. (2017) [75] presented a system for realizing Streaming SNN (S2NN) featuring switching off the parts of the implementation that are not needed. The design flow and the architecture assumed the implementation of FF-FC networks, layers of which could be ”switched off” when not in use. The design was implemented on a Xilinx Zynq-7000 device, where the integrated ARM A9 core controlled the power management. Kousanakis et al. (2017) [76] presented a system with LIF neurons on a multi-FPGA system consisting of four Xilinx Virtex-6 devices. The architecture supported learning with a BCM learning rule, as well as random synaptic current generation. The system consisted of 40 PEs responsible for calculating six neurons each, with interconnectivity information and background synaptic noise stored in external memory. Humaidi et al. (2018) [77] compared an ANN and an SNN implementations for letter recognition on FPGA. The SNN implementation supported STDP for learning the patterns. The benchmark consisted of four different letters (5x3 pixels). The SNN was reported as more area-efficient and faster than ANN. The topology used was deduced to be a 48-neuron (IZH) WTA network.
@@
Modern implementations. Ahn et al. (2020) [78] presented an architecture for realizing networks of HH neurons. The presented implementation simulated 12 million HH neurons with eight. The communication depended on the data being propagated along the axon bus from and to the PEs. The system stored duplicated information about the network state in every PE. The authors chose a neuron timestep of 0.04 ms, and the system achieved the frequency of 300 MHz on Xilinx Virtex UltraScale+. The number representation was a single-precision floating point. Ogaki et al. (2021) [79] presented an approach to realize HH-based SNNs with a fully dataflow-based approach [^2] Using the similar platform as in [80] and [81]).. The entire system was fully pipelined. The factor differentiating this architecture from the others was an explicit focus on representing the locality of biological neuronal connections - the neurons were connected only to their neighbors in the grid (2D array topology). They showed that their architecture running a 10000-neuron system (at 300 MHz) on an Alveo U200 FPGA could be up-to 7.5x faster than a state-of-the-art CPU. Zhou et al. (2021) [82] presented an NMA, which relied on a set of TDM PEs capable of simulating 1024 neurons each. They utilized a weight-sharing concept, where 16 16b weights were shared throughout the synapses of the network. The architecture supported a LIF neuron model with a linear decay, instead of exponential. The system was implemented on a Xilinx Virtex-6 device and tested with the MNIST dataset with 64 PEs to realize an FF-FC network (2010 neurons), achieving an accuracy of 98.41% after 10 timesteps.
3.2 Class 1 (Traits: fully parallel)
Figure 4: Simplified diagram of a Class 1 NMA.
@@ 3.2 Class 1 (Traits: fully parallel)
Early implementations. Maya et al. (2000) [83] presented an architecture that comprised multiple PEs that computed a pulsed neuron - three in total - to solve the XOR problem. The weights were obtained through BP in MATLAB, and the structure was clocked at 90 MHz on Virtex VX50-6. Xicotencatl et al. (2003) [84] presented a network of 1120 pulsed neurons with 1000 synapses. The architecture was implemented on Virtex-V2000 and was clocked at 35.6 MHz. The topology of the implemented network was a 2D array of neurons.
Early modern implementations. Moctezuma et al. (2013) [85] presented an architecture for simulating ion-channel dynamics and membrane voltage changes of neurons defined by HH and Traub neuron models. The models utilized floating-point arithmetic, and exponential functions were implemented as LUTs. The system was using a MicroBlaze processor for configuration and supervision, as well as communication between subunits. The entire system operated at 100 MHz and allowed the simulation of six different configurations of soma, dendrite and synapse types from the aforementioned neuron models.
3.3 Class 2 (Traits: collocated computation and memory)
@@
Early implementations. Schrauwen et al. (2008) [86] presented an LSM-based recognition system for spoken digits. The system relied on serialized synapse processing and serialized arithmetic. Architecture-wise, the system consisted of a set of PEs driven by a controller block. The neuronal dynamics of an SRM model were computed in a time-multiplexed manner, but we deduced that the PEs were not pipelined. The PEs operated at 100 MHz on Xilinx Virtex-4. The authors implemented a network of 1600 neurons as a use case, to process the incoming 16 kHz audio signal with five PEs.
Figure 5: Simplified diagram of a Class 2 NMA.
@@
Thomas et al. (2009) [87] presented an NMA using the IZH neuron model. The system allowed for simulation of 1024 neurons and 1024000 synapses, running at 100x real-time speed on a Xilinx Virtex-5 device clocked at 133 MHz. The system achieved a throughput of 2.26 GFLOPs with double-precision floating-point number representation. Architecture-wise, the system consisted of a synaptic unit, implementing four synapse pipelines, an accumulator and synaptic weight RAMs and a neuron unit with a single neuron update pipeline. Wildie et al. (2009) [88] presented an NMA that included a gap junction model on top of classic electrochemical synapses. They chose IZH as a neuron model for this system. Architecture-wise, the system consisted of eight PEs and could simulate a network of 1000 randomly-connected neurons. Every PE consisted of three pipelines - for neuron update, spike propagation and gap junctions dynamics. The system supported 24-bit fixed-point numerical representation for the parameters and conductance delay. It was clocked at 20 MHz and achieved 5.8x speed-up over real-time operation.